10 Feb 2012admin correction● Neutral
Re-lodged version of the 10 February coal seam gas conference presentation, labelled a correction. The visible text matches the original and the change is not stated.
Past year: ▲ 0%(0.60c – 0.60c AUD)
Each one in plain English, with the numbers worked through for a shareholder.
10 Feb 2012admin correction● Neutral
Re-lodged version of the 10 February coal seam gas conference presentation, labelled a correction. The visible text matches the original and the change is not stated.
10 Feb 2012investor presentation● Neutral
Conference presentation by an Icon representative, as the Managing Director was absent, on gas market trends, CSG and shale gas opportunities and shale versus CSG costs. General industry commentary; the visible text contains no new company results.
31 Jan 2012other● Neutral
Appendix 5B for December 2011 quarter: cash fell to $14.049m from $14.268m, with no financing inflows. Estimated outflow next quarter is $6.847m. Shares unchanged at 469,301,394. Routine filing.
28 Dec 2011drilling result▼ BadMajor
North Maroon DW1 tests at 1,835-1,860m and 1,568-1,593m flowed formation water, with some gas in the upper test, so no oil flowed. Icon says the sandstone is good reservoir with oil shows, but a trap is not yet defined. South Maroon and other locations are being considered.
21 Dec 2011drilling progress report● Neutral
Logs on North Maroon DW1 are evaluated and a first drill stem test is under way over 1,840-1,865m, where strong oil shows were seen. Further tests on other zones and coal seams depend on that first result.
20 Dec 2011drilling progress report● Neutral
Logging continues on North Maroon DW1 to evaluate the Precipice shows, before deciding on drill stem tests. A final decision on plugging back and sidetracking west awaits the test results. Coal seam gas sample tests are continuing.
19 Dec 2011drilling result● Neutral
North Maroon DW1 reached 1,891m. Icon says oil shows prove a hydrocarbon system in ATP 626P, but the main lower Precipice sandstone was absent here. The well may be plugged back and sidetracked west. Tests are still to come.
16 Dec 2011drilling progress report● Neutral
North Maroon DW1 is at 1,742m in the Evergreen Formation, drilling slowly through small faults. More strong fluorescent cut shows and gas of 10 units of C1 were seen at 1,731-1,736m. Icon says significance is not yet established.
15 Dec 2011drilling progress report● Neutral
North Maroon DW1 is drilling at 1,618m with weak to strong bluish-white fluorescent shows, usually a sign of residual oil, but no rise in background gas. Icon calls the shows encouraging but says their significance is unproven. Precipice target is still to come.
13 Dec 2011drilling progress report● Neutral
North Maroon DW1 reached 1,473m measured depth, passing through Walloon Coal Measures as planned. Drilling paused to fix rig and downhole equipment problems, with restart expected 14 December. Target remains about 2,100m; Icon holds 100%.
6 Dec 2011drilling progress report● Neutral
North Maroon DW1 is at 272m with surface casing cemented and blowout preventers tested, awaiting directional drilling tools. Target depth is about 2,100m, 21 days from spud. The release gives the spud date as 1 December, which conflicts with the 2 December spud announcement.
2 Dec 2011drilling spud● NeutralMajor
North Maroon DW1 was spudded in ATP 626P and drilling ahead. It targets the Precipice Sandstone and is planned to take 21 days to about 2,100m. Icon holds 100% after Stanwell's Goondi declined to participate. Spud slipped from the planned 24 November.
1 Dec 2011director officer appointment● Neutral
Chief Financial Officer Dr Kevin Jih was appointed an executive director, giving the board two executive and four non-executive directors. No extra pay beyond his existing contract. The chairman links it to the company's focus on China.
28 Nov 2011director officer appointment● NeutralMajor
Icon re-appointed Managing Director Ray James for a further five years from 31 December 2011. The chairman says this supports the strategy of establishing reserves to satisfy the gas sale agreements.
28 Nov 2011investor presentation● Neutral
AGM presentation by the Managing Director restates the strategy: shale and coal seam gas exploration in the Cooper, Surat and Gippsland basins to find reserves for the Shantou LNG deal (40 Mt over 20 years). It shows North Maroon DW1 and Tiger-1 targets. Repeats known facts.
23 Nov 2011drilling scheduled● NeutralMajor
Icon signed an Atlas Drilling contract to drill North Maroon DW1 in ATP 626P, targeting the Precipice Sandstone that produces at the Moonie oilfield. Stanwell's Goondi chose not to participate, so Icon holds 100%. A change from the coal seam gas focus; spud planned 24 November.
28 Oct 2011other▼ Bad
Appendix 5B for the September 2011 quarter: cash fell to $14.268m from $18.379m, with no financing inflows. Estimated outflow next quarter is $6.168m. Shares unchanged at 469,301,394. Routine filing.
27 Oct 2011annual report● Neutral
2011 Annual Report restates that Icon must find 2P reserves for its two sales deals: 337.5 PJ for Stanwell and 2 TCF (2,121 PJ) certified by March 2013 for the Shantou LNG deal. It estimates ATP 855P at some 20 TCF recoverable, with drilling expected in June 2012.
14 Sept 2011director officer appointment▼ Bad
Chief Operating Officer Larry Brown resigned immediately for personal reasons. Managing Director Ray James will seek a consulting engineer to cover the role until a replacement is found.
31 Aug 2011annual report▼ Bad
Full-year accounts to 30 June 2011 show a loss of $6.0m (prior year $3.6m). They record the 40% ATP 855P transfer to Beach, backdated to 5 October 2010, and the Stanwell 50% earn-in. Covering letter only; the figures are in the attached accounts.
30 Aug 2011drilling scheduled▲ GoodMajor
Icon says Beach has secured the Ensign #65 rig for the ATP 855P farm-in well, arriving April 2012 and drilling by June 2012 at the latest. Beach funds Icon's share of the ~$16m programme except Icon's $1.75m. First ATP 855P drilling timetable since the July settlement.
29 Aug 2011joint venture update▲ Good
Beach has secured two rigs for its 2012 Cooper Basin shale gas program. Ensign#65, due about April 2012, will drill the first horizontal well in ATP 855P (Icon 40%, Beach 40%). This is Beach's release; Icon's own version follows.
29 July 2011other● Neutral
Appendix 5B cash flow statement for the June 2011 quarter: cash ended at $18.379m, up from $17.938m. Next-quarter outflow is estimated at $5.39m. Shares on issue are unchanged at 469,301,394. Routine filing accompanying the quarterly report.
15 July 2011joint venture update● Neutral
Beach's own release of the 15 July settlement repeats the terms: court case discontinued, 40% of ATP 855P to Beach, and Icon takes 33.333% of the PEL 218 post-Permian joint venture. Duplicates Icon's announcement.
15 July 2011farm in farm out agreement● NeutralMajor
Icon and Beach settled the ATP 855P dispute with a signed farm-in. Icon transfers 40% to Beach (subject to Ministerial approval), leaving Icon 40%. Beach funds Icon's share of an estimated $16m work programme, except $1.75m Icon pays. Icon gets 33.333% of PEL 218 post-Permian.
27 June 2011other● Neutral
June shareholder newsletter covers the shale gas outlook for ATP 855P, NSAI certification work, a planned seismic programme and the Beach dispute (trial 18 July). Promotional update; no new announced results.
21 June 2011drilling result▼ BadMajor
Lydia-13 was plugged and abandoned at 825m because coal permeability was too low for economic gas production. That ends the four-well programme with no wells kept. A follow-up programme from 1 July, possibly outside the farm-in area, needs approvals.
14 June 2011drilling progress report● Neutral
Lydia-13 is being cored at 759m, with logging and testing to follow. It is the last well of the four-well programme this budget year. A further programme from 1 July is possible, subject to board and partner approval.
7 June 2011drilling result▼ Bad
Lydia-12 reached 900m with 8.9m of coal, and all three drill stem tests showed low permeability. The joint venture plugged and abandoned it. The rig moved to Lydia-13, spudded 6 June. Third consecutive poor result.
31 May 2011drilling progress report● Neutral
The rig returned to Lydia-12, and the stuck drilling assembly was recovered on 30 May. Drilling restarted on 31 May. (The release says it returned on 25 March, which appears to be a typo for May.)